If ethanol has a density of 0.785g/cm3, calculate the volume of 82.5g of ethanol.
1
Expert's answer
2018-09-17T08:51:18-0400

If ethanol has a density of 0.785g/cm3, calculate the volume of 82.5g of ethanol.
p=0.785 g/cm3
m= 82.5 g
V-?
Solution:
m=p*V, V=m/p= 82.5/0.785=105 cm3
Answer: V= 105 cm3